### v1

@@ Abstract
With the evolution of large language models, traditional Transformer models become computationally demanding for lengthy sequences due to the quadratic growth in computation with respect to the sequence length. Mamba, emerging as a groundbreaking architecture in the field of generative AI, demonstrates remarkable proficiency in handling elongated sequences with reduced computational and memory complexity. Nevertheless, the existing training framework of Mamba presents inefficiency with variable-length sequence inputs. Either single-sequence training results in low GPU utilization, or batched processing of variable-length sequences to a maximum length incurs considerable memory and computational overhead. To address this problem, we analyze the performance of bottleneck operators in Mamba under diverse tensor shapes and proposed PackMamba, a high-throughput Mamba that efficiently handles variable-length sequences. Diving deep into state-space models (SSMs), we modify the parallel operators to avoid passing information between individual sequences while maintaining high performance. Leveraging hardware-software co-optimization, this modification ensures coalesced memory access to position indices without extra kernel overhead. Experimental results on an NVIDIA A100 GPU demonstrate throughput exceeding the baseline single-sequence processing scheme: 3.06x speedup on the 1.4B model and 2.62x on the 2.8B model.
@@ 1 Introduction
Mamba’s training faces challenges with variable-length sequence inputs. Profiling shows that in single-sequence training, GPU tasks are fine-grained, with large gaps between tasks. These gaps, caused by frequent CPU-GPU synchronization and high kernel launch overhead, lead to inefficient GPU utilization. Anothor approach is pad to maximum length for batch training. However, these padded zeros introduce significant computation and memory overhead.
@@
To address the issue of low training throughput in Mamba, we introduce PackMamba - a high-performance Mamba that supports variable-length sequences. As shown in Figure 1, PackMamba initially packs variable-length sequences into longer sequences(4096 in Mamba-1.3B) using a indices to record the original sequence start positions within the long sequence. By modifying the selective scan and convolution operators to prevent the cross-sequence access(marked in red line), the training process is adapted to support variable-length sequences. Additionally, we employ techniques such as shared memory and memory access merging for a synergistic acceleration of GPU memory access, optimizing performance. These advancements ensure that PackMamba not only overcomes the limitations of Mamba but also significantly enhances its efficiency and scalability, thereby making it a more robust solution for neural network training.
@@
    We conduct an operator analysis of the SSM, the bottleneck of Mamba, and discovered that hardware and the model demonstrate varying affinities towards sequence length.
@@ 2.1 Variable-Length Inputs Train
As depicted in Section 1, a straightforward solution to solve fine-grained computation in variable-length training is to pad all sequences with zeros to the maximum sequence length, but this method leads to a significant waste of memory and computational resources. In our experiment, it has been demonstrated that the use of this padding method with the InternLM dataset [10, 2] results in 66.3% padding rate. As shown in Figure 6 - baseline, indicate that the SSM operator has become a performance bottleneck and involves a significant amount of idle computation,
There has been some work focusing on the training issue of variable-length sequences in Transformers. Both Tencent [4] and Baidu [13] initially attempted to minimize the proportion of zero-padding by bundling sequences of similar lengths. However, this approach doesn’t completely solve the problem. ByteTransformer [14] proposed a padding-free algorithm that packs the input tensor with variable-length sequences and calculates the positioning offset vector for all transformer operations to index. This approach, combined with software and hardware co-optimization, achieved good results. In addition, ByteTransformer also employs methods such as kernel fusion and CUTLASS grouped GEMM optimizations to enhance overall performance.
@@ 2.2 SSM Operator Analysis
In the padding approach, the SSM operator is a bottleneck, using 59.3% of the process resources. zero-padding along the sequence dimension wastes its computational resources. Therefore, we analyze the performance of the operator with different seqlen inputs. Analyzing Figure 2 and the details of the operator, we draw the following conclusions:
@@
    When $2^{n}<\text{seqlen}<2^{n+1}$, the duration increases slowly with the increase in seqlen. This is due to the operator’s internal padding zeros.
@@
    When $\text{seqlen}=2^{n}$ (or a multiple of 2048), the processing time significantly decreases. This reduction is due to an internal vector loading mechanism that can accelerate processing by about $1.51$ to $2.03$ times when activated.
@@
Therefore, constructing input sequence lengths as input$(\text{seqlen}=2^{n})$ can achieve a high throughput. As for the external padding of zeros, one method is to pack multiple sequences into one long set, but passing states between sequences can lead to contamination of input data. We have analyzed and modified Mamba to ensure consistent training results before and after packing, thereby accelerating training.
@@ 3.1 Packing-Unpacking Invariance
The pack() operation refers to concatenating the input tensor along the sequence dimension to obtain a packed_sequence and the auxiliary structure position_indices (detailed in Section Section 3.3) as shown in Figure 3(a). The unpack() operation is the inverse of pack().
@@ 3.4 SSM_pack Implementation
Modifying the input of scanMul to set $\overline{A}_{position\_indices=0}\to 0$ ensures this forward parallel operations with PUI too. This is because, if the $a^{th}$ term is set to zero, for the $m^{th}$ item processed by a single thread in scanMul: before the step, if the multiplied term $n>a$, then $A[m]\neq 0$; if $n<a$, then $A[m]=0$.
@@
Similarly, the backward process consists of another two scan operator, where modifications only require setting $\overline{A}_{position\_indices=0}\to 0$
@@ 4 Evaluation
We evaluate our approach on an NVIDIA A100 GPUs. The training data, sourced from authentic corpora, consisted of sequences ranging in length from 57 to 2048, with an average length of 646. We utilize bfloat16 and float32 tensor types for inputs. The models trained were Mamba-110m with 16 layers and 1024 dimensions, Mamba-1.4B with 48 layers and 2048 dimensions, and Mamba-2.8B with 64 layers and 2560 dimensions. Training was executed using an 8-GPU data parallel. We compare three approaches: single-sequence, padding, and pack.
@@
Kernel Speedup Analysis. We switch the baseline to the padding approach, as the kernel duration of the single-sequence approach is too sparse for reliable comparison. As shown in Figure 5, the forward-backward process achieves a 3.91x speedup, primarily due to the reduction in GEMM and SSM kernel durations. The increased density of the packed sequences eliminates significant amounts of idle computations, showcasing a smaller acceleration space for memory-intensive operations like conv1d.
@@ 5 Discussion
PackMamba supports packing training with variable-length sequences while ensuring mathematical equivalence. Moreover, there are no instances of sequences spanning across packed sequences. In future work, we plan to address this issue by allowing sequences to be cut into two parts at the end of long sequences, with states still being passed between these parts. This approach will reduce padding to zero while maintaining data integrity and even support parallel strategies for infinitely long sequences.

### v2

@@ Abstract
With the evolution of large language models, traditional Transformer models become computationally demanding for lengthy sequences due to the quadratic growth in computation with respect to the sequence length. Mamba, emerging as a groundbreaking architecture in the field of generative AI, demonstrates remarkable proficiency in handling elongated sequences with reduced computational and memory complexity. Nevertheless, the existing training framework of Mamba presents inefficiency with variable-length sequence inputs. Either single-sequence training results in low GPU utilization, or batched processing of variable-length sequences to a maximum length incurs considerable memory and computational overhead. To address this problem, we analyze the performance of bottleneck operators in Mamba under diverse tensor shapes and propose PackMamba, a high-throughput Mamba that efficiently handles variable-length sequences. Diving deep into state-space models (SSMs), we modify the parallel operators to avoid passing information between individual sequences while maintaining high performance. Leveraging hardware-software co-optimization, this modification ensures coalesced memory access to position indices without extra kernel overhead. Experimental results on an NVIDIA A100 GPU demonstrate throughput exceeding the baseline single-sequence processing scheme: 3.06x speedup on the 1.4B model and 2.62x on the 2.8B model.
@@ 1 Introduction
Mamba’s training faces challenges with variable-length sequence inputs. Profiling shows that in single-sequence training, GPU tasks are fine-grained, with large gaps between tasks. These gaps, caused by frequent CPU-GPU synchronization and high kernel launch overhead, lead to inefficient GPU utilization. Another approach is pad to maximum length for batch training. However, these padded zeros introduce significant computation and memory overhead.
@@
To address the issue of low training throughput in Mamba, we introduce PackMamba - a high-performance Mamba that supports variable-length sequences. As shown in Figure 1, PackMamba initially packs variable-length sequences into longer sequences (4096 in Mamba-1.3B) using indices to record the original sequence start positions within the long sequence. By modifying the selective scan and convolution operators to prevent the cross-sequence access (marked in red line), the training process is adapted to support variable-length sequences. Additionally, we employ techniques such as shared memory and memory access merging for a synergistic acceleration of GPU memory access, optimizing performance. These advancements ensure that PackMamba not only overcomes the limitations of Mamba but also significantly enhances its efficiency and scalability, thereby making it a more robust solution for neural network training.
@@
    We conduct an operator analysis of the SSM, the bottleneck of Mamba, and discover that hardware and the model demonstrate varying affinities towards sequence length.
@@ 2.1 Variable-Length Inputs Train
As depicted in Section 1, a straightforward solution to solve fine-grained computation in variable-length training is to pad all sequences with zeros to the maximum sequence length, but this method leads to a significant waste of memory and computational resources. Our experiments demonstrate that the use of this padding method with the InternLM dataset [10, 2] results in 66.3% padding rate.
There has been some work focusing on the training issue of variable-length sequences in Transformers. Both Tencent [4] and Baidu [13] initially attempted to minimize the proportion of zero-padding by bundling sequences of similar lengths. ByteTransformer [14] proposed a padding-free algorithm that packs the input tensor with variable-length sequences and calculates the positioning offset vector for all transformer operations to index. This approach, combined with software and hardware co-optimization, achieved good results. In addition, ByteTransformer also employs methods such as kernel fusion and CUTLASS grouped GEMM optimizations to enhance overall performance.
@@ 2.2 SSM Operator Analysis
As shown in Figure 6, in the padding approach, the SSM operator is a bottleneck, using 59.3% of the process resources. Zero-padding along the sequence dimension wastes its computational resources. Therefore, we analyze the performance of the operator with different seqlen inputs. Analyzing Figure 2 and the details of the operator, we draw the following conclusions:
@@
    When $2^{n}<\text{seqlen}<2^{n+1}$, the duration increases slowly with increasing seqlen. This is due to the operator’s internal padding zeros.
@@
    When $\text{seqlen}=2^{n}$ (or a multiple of 2048), the processing time significantly decreases. This reduction is due to an internal vector loading mechanism which provides a speedup of 1.51x to 2.03x when activated.
@@
Therefore, constructing input sequences such that $\text{seqlen}=2^{n}$ can achieve high throughput. As for the external padding of zeros, one method is to pack multiple sequences into one long set, but passing states between sequences can lead to contamination of input data. We have analyzed and modified Mamba to ensure consistent training results before and after packing, thereby accelerating training.
@@ 3.1 Packing-Unpacking Invariance
The pack() operation refers to concatenating the input tensor along the sequence dimension to obtain a packed_sequence and the auxiliary structure position_indices (detailed in Section 3.3) as shown in Figure 3(a). The unpack() operation is the inverse of pack().
@@ 3.4 SSM_pack Implementation
Modifying the input of scanMul to set $\overline{A}_{position\_indices=0}\to 0$ ensures this forward parallel operation with PUI too. This is because, if the $a^{th}$ term is set to zero, for the $m^{th}$ item processed by a single thread in scanMul: before the step, if the multiplied term $n>a$, then $\overline{A}[m]\neq 0$; if $n<a$, then $\overline{A}[m]=0$.
@@
Similarly, the backward process consists of another two scan operators, where modifications only require setting $\overline{A}_{position\_indices=0}\to 0$
@@ 4 Evaluation
We evaluate our approach on NVIDIA A100 GPUs. The training data, sourced from authentic corpora, consisted of sequences ranging in length from 57 to 2048, with an average length of 646. We utilize bfloat16 and float32 tensor types for inputs. The models trained were Mamba-110m with 16 layers and 1024 dimensions, Mamba-1.4B with 48 layers and 2048 dimensions, and Mamba-2.8B with 64 layers and 2560 dimensions. Training was executed using an 8-GPU data parallel. We compare three approaches: single-sequence, padding, and pack.
@@
Kernel Speedup Analysis. We switch the baseline to the padding approach, as the kernel duration of the single-sequence approach is too sparse for reliable comparison. As shown in Figure 5, the forward-backward process achieves a 3.91x speedup, primarily due to the reduction in GEMM and SSM kernel durations. The increased density of the packed sequences eliminates significant amounts of idle computations.
@@ 5 Discussion
PackMamba supports packing training with variable-length sequences while ensuring mathematical equivalence. Moreover, there are no instances of sequences spanning across packed sequences. In future work, we plan to address this issue by allowing long sequences to be cut at the end into two parts while still passing states between these parts. This approach will reduce padding to zero while maintaining data integrity and even support parallel strategies for infinitely long sequences.